Demand is given by P = 24 – 4Q, where P is the price of the good and Q is the quantity
demanded. At P = 12, the own price (point) elasticity is
a. 0.25
b. 0.5
c. 1
d. 2
1
Expert's answer
2020-03-20T09:47:39-0400


Point elasticity is given by "\frac{\Delta{Q}}{\Delta{P}}\times{\frac{P}{Q}}"

"-4Q=P-24"

"Q=6-\frac{P}{4}"

"\frac{\Delta{Q}}{\Delta{P}}=-\frac{1}{4}"

"P=12"

"Q(12)=6-\frac{12}{4}=3"

"E=-\frac{1}{4}\times{\frac{12}{3}}=-1"

Thus, The correct answer is (c).